# HOI-Dyn: Learning Interaction Dynamics for Human-Object Motion Diffusion

### v2

@@ Abstract
Generating realistic 3D human-object interactions (HOIs) remains a challenging task due to the difficulty of modeling detailed interaction dynamics. Existing methods treat human and object motions independently, resulting in physically implausible and causally inconsistent behaviors. In this work, we present HOI-Dyn, a novel framework that formulates HOI generation as a driver-responder system, where human actions drive object responses. At the core of our method is a lightweight transformer-based interaction dynamics model that explicitly predicts how objects should react to human motion. To further enforce consistency, we introduce a residual-based dynamics loss that mitigates the impact of dynamics prediction errors and prevents misleading optimization signals. The dynamics model is used only during training, preserving inference efficiency. Through extensive qualitative and quantitative experiments, we demonstrate that our approach not only enhances the quality of HOI generation but also establishes a feasible metric for evaluating the quality of generated interactions.
@@ 2.1 Driver-Responder Synchronization
Driver-Responder Synchronization (master-slave synchronization) is observed in many coupled systems, such as those found in biological and physical domains [17, 18, 19]. These systems often rely on controllers to achieve synchronization, employing strategies such as adaptive and feedback control [20]. From a control theory perspective, synchronization is achieved by adjusting the coupling strength and applying feedback to align the system states [21]. This concept has been applied to various domains, but similar challenges arise in HOI generation.
Current HOI generation frameworks, however, often treat human and object motions as independent entities [13, 22, 23], which can lead to physically implausible object behaviors. Existing generative approaches [24] model the overall consistency between human and object, but they fail to account for the fine-grained interaction dynamics at the point of contact. To address this gap, we propose conceptualizing HOI generation as a Driver-Responder System, where human actions serve as the driver, controlling the object’s response. This perspective ensures a more coherent and physically consistent interaction by capturing the causal relationship between human movements and object reactions. We argue that this approach overcomes the limitations of previous models by introducing a refined internal control mechanism that improves the physical realism and causal consistency of the generated interactions.
2.2 Object and Contact Modeling
A substantial body of work on HOI generation has been dedicated to modeling object affordances and contact prediction. The method HOI-Diff [25] introduces an affordance prediction module to constrain interaction regions, by incorporating the estimated contact points into classifier-guidance to encourage close proximity between humans and objects. This method is further enhanced in CG-HOI [15] by explicitly predicting contact fields on the human body mesh and using them as strong priors during both training and inference. The method OMOMO [1] adopts a two-stage pipeline: it first infers contact points and then generates HOI sequences conditioned on these points. While these methods successfully model static interaction cues, they largely overlook the underlying interaction dynamics, i.e. how objects should physically and causally respond to human actions over time. In practice, accurate contact prediction is inherently difficult, and even with correct contact estimation, these approaches often fail to produce temporally coherent and physically realistic object behaviors.
2.3 Joint HOI Motion Generation
Another line of research explores joint generation of human-object motion using diffusion-based models. The model CHOIS [13] encodes human poses, object states, and limb-contact labels as part of the noising-denoising process. Their work enforces contact constraints by introducing additional object pose supervision during training and applying hard constraints during sampling. The method THOR [14] leverages a relationally-aware diffusion model that initiates human-object interaction guided by text, and intervenes in object motion based on learned human-object relations at each diffusion step.
While these methods integrate human and object motions within a unified generative framework and produce globally coherent sequences, they typically treat interaction as a high-level constraint or relational prior. Consequently, they often fail to model fine-grained interaction dynamics—specifically, how an object should physically and causally respond to human motion. This implicit treatment leads to plausible-looking results that may still violate physical consistency or lack local causal coherence.
@@ 3.1 Interaction Dynamics
In practice, we observe that existing diffusion models often fail to generate causally consistent interactions due to the absence of precise control signal $u^{(t)}$, which is essential for aligning the object’s motion with human intent [13, 14]. To address this limitation, we propose the interaction dynamics as a supervisory signal to implicitly optimize it in the diffusion process. Based on  (1), we derive the object’s relative motion as
@@
The above predicted motion is further represented as a rigid-body transformation comprising a rotation ^$\hat{\mathcal{R}}^{(t\to t+k)}\in\mathrm{SO}(3)$ and translation $\hat{\mathcal{T}}^{(t\to t+k)}\in\mathbb{R}^{3}$, i.e., ${\Delta o}^{*}_{=}[\hat{\mathcal{R}}|\hat{\mathcal{T}}]$, applied to a set of object’s points $\mathcal{P}^{(t)}$ to yield the predicted future configuration as follows:
@@
To ensure that ^$\hat{\mathcal{R}}^{(t\to t+k)}$ is a valid rotation matrix, we apply singular value decomposition (SVD)-based projection to the raw network output ~$\tilde{\mathcal{R}}\in\mathbb{R}^{3\times 3}$, such that ~$\tilde{\mathcal{R}}=U\Sigma V^{\top}$ and ^$\hat{\mathcal{R}}=UV^{\top}$. We then define the following cost for object dynamics to quantify the error between the transformed keypoints:
@@ 3.2 Conditional Motion Diffusion
Although the standard loss function (9) yields plausible motion sequences, physical inconsistencies—particularly in object trajectories—persist. To address this, we introduce an auxiliary Interaction Dynamics Loss based on a shared dynamics model $\mathcal{D}$. This loss penalizes discrepancies in object motion between generated and ground-truth sequences as follows:
@@
To mitigate the influence of these errors, we assume that $\mathcal{D}$ is locally smooth and time-homogeneous—its bias depends on the state but not on time $t$. This is reasonable since $\mathcal{D}$ predicts object dynamics from current states without explicit time dependence, and local smoothness reflects the continuity of physical motion. Under this assumption, when the generative and ground-truth trajectories exhibit similar state distributions, the error residuals tend to cancel out in expectation.
@@
This residual formulation provides a robust supervisory signal: even if $\mathcal{D}$ is imperfect, its bias is removed through subtraction, allowing the learning algorithm to focus on the true inconsistencies in the generated motion. For a more detailed discussion, see Appendix C.
With this loss in place, each step of the generation process is guided to yield a more accurate control signal $u_{t}$, resulting in physically grounded and temporally coherent HOI sequences. We therefore extend the standard HOI training objective $\mathcal{L}_{\text{hoi}}$ (9) by incorporating both $\mathcal{L}_{\text{dyn}}$ (10) and an object-level reconstruction loss $\mathcal{L}_{\text{obj}}=\mathbb{E}_{t}[\Phi(o_{t},\hat{o}_{t})]$. All these losses are equally weighted to simplify the overall training process and avoid the need for complex hyperparameter tuning.
@@ 4 Experiments
We train and evaluate HOI generation using two datasets: (i) FullBodyManipulation, which provides 10 hours of high-quality paired object and human motion data involving 15 different objects [1]. We apply the OMOMO partitioning to this dataset: 15 subjects for training and 2 for testing. (ii) 3D-FUTURE, which consists of 3D models of various furniture items [29]. To assess the generalization capability, we select 17 objects and pair them with motion sequences from the FullBodyManipulation testing set.
@@
We train the interaction dynamics using FullBodyManipulation with $K=2$ as the maximum prediction horizon. The network has 0.5M parameters. We use the Adam optimizer with a learning rate of $1\times 10^{-3}$, adjusted via CosineAnnealingWarmRestarts [31] for 150 epochs with a batch size of 32. The model is then transferred to the HOI motion diffusion task, where fine-grained driver-responder relationships are encouraged during denoising, as in (10). Training is done from scratch with a learning rate of $1\times 10^{-4}$, batch size 32, and 100,000 steps. All experiments are conducted on a single NVIDIA RTX A4500 GPU, with total training time around 10 hours.
@@ 4.1 Qualitative Results
As shown in Figs. 2 (a) and (b), CHOIS, which lacks interaction dynamics modeling, often causes premature object motion. The objects tend to move toward expected contact points (e.g., the hand) spontaneously before human actions begin, leading to implausible behaviors such as bouncing, wobbling, or sliding. In contrast, HOI-Dyn reduces such artifacts, promoting more natural human-object interactions. Fig. 2 (c) further highlights the difference in object responses after contact: HOI-Dyn generates subtle, physically plausible reactions (with slight sliding after a gentle kick), while CHOIS often produces exaggerated effects like flying or floating.
Fig. 2 (d) shows the keyframes for sequence-level comparison. Both methods generally respect input constraints, demonstrating the effectiveness of motion diffusion. However, CHOIS often exhibits disjoint human-object motion and weak causality, while HOI-Dyn ensures consistent object motion and maintains contact, resulting in more coherent and plausible HOI sequences. More visualization results can be found in AppendixD.
@@ 4.2 Quantitative Results
[…]
@@
[…]
@@ 4.3 Further Discussion
We analyze the effects of prediction horizon $K$, network design, and model complexity on the interaction dynamics performance. As shown in Fig. 4, a small $K$ limits large motion capture, while a large $K$ weakens subtle interaction modeling. Empirically, $K{=}2$ or $K{=}3$ yield the best results. For network design, we compare the decoupled and coupled motion strategies. Inspired by [34, 35, 36], the decoupled approach predicts object rotation and translation separately. Our coupled design instead models the unified influence of human motion and contact, yielding a better performance under similar parameter and FLOP constraints, highlighting the inherent coupling in HOI, as shown in Table 3. We also evaluate the model efficiency by varying the Transformer depth ($D$), feature dimension ($F$), and head count ($H$). The results in Table 3 show that a lightweight model with 0.5M parameters and 0.2 GFLOPs suffices to capture high-quality interaction dynamics. For a more detailed discussion, see Appendix E
Effect of Different Guidance.
Classifier-based guidance is often employed during inference in generative models. Following [13], we apply two types of guidance: feet-floor and hand-object. For a more detailed description, see Appendix F. As shown in Table 4, even without guidance, our method surpasses CHOIS, especially in the interaction quality. The feet-floor term enhances foot realism and reduces FID, while the hand-object term boosts contact accuracy but slightly harms motion quality. Combining both yields consistent gains across all metrics, achieving a new SOTA.
Dynamics as Causality Metric.
@@
Decoupled (K=2) (D1-F64-H8)×$\times 2$ 0.358±0.0006 0.532±0.0009 0.714±0.0027 0.919±0.0037 1.144±0.0057 2.613±0.0059 0.463 0.108
Decoupled (K=2) (D2-F64-H8)×$\times 2$ 0.340±0.0011 0.503±0.0006 0.676±0.0007 0.870±0.0015 1.080±0.0071 2.537±0.0049 0.496 0.200
Decoupled (K=2) (D4-F64-H8)×$\times 2$ 0.337±0.0006 0.503±0.0004 0.676±0.0034 0.866±0.0022 1.062±0.0047 2.409±0.0049 0.564 0.385
@@
[…]
@@
Figure 5: Frame-wise object point cloud loss. Our method HOI-Dyn (green line), consistently achieves lower loss than baselines without dynamics supervision (blue line). Notably, high loss peaks correspond to physically implausible HOI cases, which are effectively identified by our dynamics model.
5 Limitation and Conclusion

### v3

@@ Abstract
Generating realistic 3D human-object interactions (HOIs) remains a challenging task due to the difficulty of modeling detailed interaction dynamics. Existing methods treat human and object motions independently, resulting in physically implausible and causally inconsistent behaviors. In this work, we present HOI-Dyn, a novel framework that formulates HOI generation as a driver-responder system, where human actions drive object responses. At the core of our method is a lightweight transformer-based interaction dynamics model that explicitly predicts how objects should react to human motion. To further enforce consistency, we introduce a residual-based dynamics loss that mitigates the impact of dynamics prediction errors and prevents misleading optimization signals. The dynamics model is used only during training, preserving inference efficiency. Through extensive qualitative and quantitative experiments, we demonstrate that our approach not only enhances the quality of HOI generation but also establishes a feasible metric for evaluating the quality of generated interactions. Project website:https://wulin97.github.io/hoi-dyn
@@ 2 Related Work
2.1 Object-Guided HOI Generation
A subset of HOI generation methods leverages object trajectories or waypoints to guide human motion. OMOMO [1] takes a full sequence of object states as input and generates the corresponding human poses, whereas CHOIS [13] and Wu et al. [2] rely on sparse object waypoints (e.g., roughly one waypoint every 30 frames), leaving the object’s detailed responses to be implicitly determined by the model. These methods typically add auxiliary supervision or constraints to encourage plausible human-object contact. However, such guidance mainly steers the diffusion process toward ensuring that contact occurs, rather than modeling the underlying interaction itself. While this strategy can yield globally coherent sequences, it remains at a high level: the model is not trained to capture how objects physically react to human actions in a fine-grained and causally consistent manner, leaving the central challenge of realistic HOI underexplored.
2.2 Joint Human-Object Motion Generation
Another line of work generates HOIs jointly without conditioning on future object states. HOI-Diff [17] relies on affordance prediction and estimated contact points to guide interactions, while CG-HOI [15] leverages contact fields on the human mesh as strong priors. THOR [14] models relational cues for coordinated motions, HIMO [18] employs a dual-branch conditional diffusion with a mutual interaction module for cross-modal fusion, and ChainHOI [19] adopts a spatiotemporal graph architecture with a kinematics-aware module to capture joint- and chain-level dependencies.
Despite these advances, common challenges remain. First, predicting accurate contact points or modeling interactions precisely at contact regions is inherently difficult. Second, although high-level constraints or interaction modules encourage consistency, they often fail to capture the causal and fine-grained dynamics of object responses to human actions. As a result, object behaviors may appear temporally inconsistent or physically implausible, even when global plausibility is achieved. This highlights a missing perspective: explicitly treating human actions as the driver and object motions as their causal responses. Such a formulation is crucial for achieving physically realistic HOI generation.
@@ 2.3 Driver-Responder Synchronization
Driver-Responder Synchronization, also referred to as master-slave synchronization, has been widely observed in coupled systems across biological and physical domains [20, 21, 22]. In these systems, controllers achieve synchronization through adaptive and feedback strategies [23, 24]. Drawing inspiration from these systems, we conceptualize HOI generation as a Driver-Responder System, where human actions serve as the driver, controlling the object’s response. Unlike prior methods that treat human and object motions independently [13, 25, 26], this perspective captures the causal relationship between human movements and object reactions, ensuring more coherent and physically consistent interactions. By introducing an internal control mechanism, the Driver-Responder formulation provides a principled framework for generating HOIs with refined temporal coherence and realistic object dynamics, addressing the limitations of previous high-level constrained approaches.
@@ 3.1 Interaction Dynamics
In practice, we observe that existing diffusion models often fail to generate causally consistent interactions due to the absence of precise control signal $u^{(t)}$, which is essential for aligning the object’s motion with human intent [13, 14]. To address this limitation, we propose the interaction dynamics as a supervisory signal to implicitly optimize it in the diffusion process. Based on  (1), we can derive the object’s relative motion as
@@
The above predicted motion is further represented as a rigid-body transformation comprising a rotation $\hat{\mathcal{R}}^{(t\to t+k)}\in\mathrm{SO}(3)$ and translation $\hat{\mathcal{T}}^{(t\to t+k)}\in\mathbb{R}^{3}$, i.e., ${\Delta o}^{*}_{=}[\hat{\mathcal{R}}|\hat{\mathcal{T}}$], applied to a set of object’s points $\mathcal{P}^{(t)}$ to yield the predicted future configuration as follows:
@@
To ensure that $\hat{\mathcal{R}}^{(t\to t+k)}$ is a valid rotation matrix, we apply singular value decomposition (SVD)-based projection to the raw network output $\tilde{\mathcal{R}}\in\mathbb{R}^{3\times 3}$, such that $\tilde{\mathcal{R}}=U\Sigma V^{\top}$ and $\hat{\mathcal{R}}=UV^{\top}$. We then define the following cost for object dynamics to quantify the error between the transformed keypoints:
@@ 3.2 Conditional Motion Diffusion
Although the standard loss function (9) produces plausible motion sequences, physical inconsistencies persist, particularly in object trajectories. To address this, we introduce an auxiliary Interaction Dynamics Loss based on a shared dynamics model $\mathcal{D}$. This loss penalizes discrepancies in object motion between generated and ground-truth sequences, enforcing a causal alignment between human actions and object responses, as follows:
@@
To mitigate the influence of these errors, we assume that $\mathcal{D}$ is locally smooth and time-homogeneous—its bias depends on the state but not on time $t$. This assumption is reasonable because $\mathcal{D}$ predicts object dynamics from current states without explicit time dependence, and local smoothness reflects the continuity of physical motion (see Appendix E.2 for a detailed discussion of these assumptions). Under this assumption, when the generative and ground-truth trajectories exhibit similar state distributions, the error residuals tend to cancel out in expectation:
@@
This residual formulation provides a robust supervisory signal: even if $\mathcal{D}$ is imperfect, its bias is removed through subtraction, allowing the learning algorithm to focus on the true inconsistencies in the generated motion. For a more detailed discussion, see Appendix C. With this loss in place, each step of the generation process is guided to yield a more accurate control signal $u_{t}$, resulting in physically grounded and temporally coherent HOI sequences. We therefore extend the standard HOI training objective $\mathcal{L}_{\text{hoi}}$ (9) by incorporating both $\mathcal{L}_{\text{dyn}}$ (10) and an object-level reconstruction loss $\mathcal{L}_{\text{obj}}=\mathbb{E}_{t}[\Phi(o_{t},\hat{o}_{t})]$. All these losses are equally weighted to simplify the overall training process and avoid the need for complex hyperparameter tuning.
@@ 4 Experiments
We train and evaluate HOI generation using two datasets: (i) FullBodyManipulation, which provides 10 hours of high-quality paired object and human motion data involving 15 different objects [1]. We apply the OMOMO partitioning to this dataset: 15 subjects for training and 2 for testing; and (ii) 3D-FUTURE, which consists of 3D models of various furniture items [30]. To assess the generalization capability, we select 17 objects and pair them with motion sequences from the FullBodyManipulation testing set.
@@
We train the interaction dynamics using FullBodyManipulation with $K=2$ as the maximum prediction horizon. The Transformer-based network has 0.5M parameters. We use the Adam optimizer [32] with a learning rate of $1\times 10^{-3}$, adjusted via CosineAnnealingWarmRestarts [33] (with $T_{0}=20$, $T_{\text{mult}}=1$, $\eta_{\text{min}}=1\times 10^{-5}$), for 150 epochs with a batch size of 32. The model is then transferred to the HOI motion diffusion task, where fine-grained driver-responder relationships are encouraged during denoising, as in (10). Training is done from scratch with a learning rate of $1\times 10^{-4}$, batch size 32, and 100,000 steps. All experiments are conducted on a single NVIDIA RTX A4500 GPU, with total training time around 10 hours.
@@ 4.1 Qualitative Results
As shown in Figs. 2 (a) and (b), CHOIS, which lacks interaction dynamics modeling, often causes premature object motion. The objects tend to move toward expected contact points (e.g., the hand) spontaneously before human actions begin, leading to implausible behaviors such as bouncing, wobbling, or sliding. In contrast, HOI-Dyn reduces such artifacts, promoting more natural interactions. Fig. 2 (c) further highlights the difference in object responses after contact: HOI-Dyn generates subtle, physically plausible reactions (with slight sliding after a gentle kick), while CHOIS often produces exaggerated effects like flying or floating. Fig. 2 (d) shows the keyframes for sequence-level comparison. Both methods generally respect input constraints, demonstrating the effectiveness of motion diffusion. However, CHOIS often exhibits disjoint human-object motion and weak causality, while HOI-Dyn ensures consistent object motion and maintains contact, resulting in more coherent and plausible HOI sequences. More visualization results can be found in AppendixD.
@@ 4.2 Quantitative Results
$T_{s}$ $\downarrow$ $T_{e}$ $\downarrow$ $T_{xy}$ $\downarrow$ $H_{\text{feet}}$ $\downarrow$ FS $\downarrow$ FID $\downarrow$ $C_{F1}$ $\uparrow$ C% $\uparrow$ $P_{\text{hand}}$ $\downarrow$ MPJPE $\downarrow$ $T_{\text{root}}$ $\downarrow$ $T_{\text{obj}}$ $\downarrow$ $R_{\text{obj}}$ $\downarrow$
@@
$T_{s}$ $\downarrow$ $T_{e}$ $\downarrow$ $T_{xy}$ $\downarrow$ $H_{\text{feet}}$ $\downarrow$ FS $\downarrow$ FID $\downarrow$ C% $\uparrow$ $P_{\text{hand}}$ $\downarrow$
@@
4.3 Application in 3D Scene
We demonstrate the practical applicability of our method by synthesizing dynamic HOI within realistic 3D environments, conditioned on text descriptions. Specifically, we use 3D scenes from the Replica dataset [36] and manually define instructions such as “pull the floor lamp and move it next to the sofa”. The pipeline involves three steps: (i) parsing the textual description to identify the intended interaction and associated object(s), (ii) specifying the target object’s initial placement within the scene, and (iii) planning a collision-free navigation trajectory using Habitat. The resulting path is then provided to our HOI generation model, which produces coherent and physically plausible motion sequences aligned with the instruction.
As illustrated in Fig. 3, our model successfully generates realistic agent behaviors in complex environments. In (a), the agent interacts with a floor lamp and repositions it near a sofa, while in (b), it moves a large box across the room. Both examples highlight the model’s ability to generate environment-conscious and interaction-consistent motions, with potential applications in animation, virtual reality, and robotics.
(a) Interaction with a floor lamp
(b) Interaction with a large box
Figure 3: HOI generation in realistic 3D scenes. The virtual agent interacts with different objects while maintaining physical plausibility and environmental consistency.
@@ 4.4 Further Discussion
We analyze the effects of prediction horizon $K$, network design, and model complexity on the interaction dynamics performance. As shown in Fig. 5, a small $K$ limits large motion capture, while a large $K$ weakens subtle interaction modeling. Empirically, K=2K{=}2 or K=3K{=}3 yield the best results. For network design, we compare the decoupled and coupled motion strategies. Inspired by [37, 38, 39], the decoupled approach predicts object rotation and translation separately. Our coupled design instead models the unified influence of human motion and contact, yielding a better performance under similar parameter and FLOP constraints, highlighting the inherent coupling in HOI, as shown in Table 3. We also evaluate the model efficiency by varying the Transformer depth ($D$), feature dimension ($F$), and head count ($H$). The results in Table 3 show that a lightweight model with 0.5M parameters and 0.2 GFLOPs suffices to capture high-quality interaction dynamics. For a more detailed discussion, see Appendix E
@@
[…]
Effect of Different Guidance.
Classifier-based guidance is often employed during inference in generative models. Following [13], we apply two types of guidance: feet-floor and hand-object. The feet-floor term encourages physical plausibility during locomotion and standing still by penalizing unnatural foot height above the ground, while the hand-object term enforces physically consistent contact and temporal coherence for hand–object interactions. For a more detailed description, see Appendix F. As shown in Table 4, even without guidance, our method surpasses CHOIS, especially in the interaction quality. The feet-floor term enhances foot realism and reduces FID, while the hand-object term boosts contact accuracy but slightly harms motion quality. Combining both yields consistent gains across all metrics, achieving a new SOTA.
@@
$T_{s}$ $\downarrow$ $T_{e}$ $\downarrow$ $T_{xy}$ $\downarrow$ $H_{\text{feet}}$ $\downarrow$ FS $\downarrow$ F ID $\downarrow$ $C_{F1}$ $\uparrow$ C% $\uparrow$ $P_{\text{hand}}$ $\downarrow$ MPJPE $\downarrow$ $T_{\text{root}}$ $\downarrow$ $T_{\text{obj}}$ $\downarrow$ $O_{\text{obj}}$ $\downarrow$
@@
Effect of Removing Object Waypoints.
To evaluate the intrinsic capability of our interaction dynamics model, we consider a without waypoint (w/o WP) setting, removing predefined object waypoints for both the Baseline (CHOIS) and HOI-Dyn models. In this setting, the GT difference metric is no longer meaningful, so we focus on metrics reflecting physical plausibility and interactive realism: foot sliding, penetration (any unrealistic interpenetration involving human, object, or floor), contact F1 and accuracy, FID, and diversity. Classifier-based guidance is not applied during inference, ensuring that results reflect the generator’s learned dynamics. As shown in Table 5, HOI-Dyn demonstrates improved foot and contact quality, better FID and diversity, and comparable penetration, highlighting the reasoning and generalizability of the model.
Table 5: Performance under the without waypoint setting. HOI-Dyn improves foot and contact quality, FID, and diversity, while preserving penetration, showing its intrinsic interaction dynamics.
Method FS $\downarrow$ Penetration ↓ $_{F1} \uparrow CC_{\%}$ $\uparrow$ FID ↓ Diversity ↑
CHOIS w/o WP 0.401 0.581 0.573 0.648 5.36 7.90
HOI-Dyn w/o WP 0.376 0.582 0.592 0.670 4.81 8.09
Dynamics as Causality Metric.
@@
Figure 6: Frame-wise object point cloud loss. Our method HOI-Dyn (green line), consistently achieves lower loss than the baseline without dynamics supervision (blue line). Notably, high loss peaks correspond to physically implausible HOI cases, which are effectively identified by our dynamics model.
5 Conclusion and Limitation
In this work, we emphasize the role of interaction dynamics in ensuring physical and causal consistency in HOI generation. As an early exploration, we propose a novel driver-responder framework that explicitly models these dynamics and integrates with existing HOI motion diffusion techniques to achieve more realistic human-object interactions. While alternative approaches based on physics simulators might seem capable of modeling object dynamics more directly [40], their black-box nature, lack of differentiability, and strong reliance on precise physical properties make them difficult to incorporate into generative frameworks. A detailed discussion is provided in Appendix C.4.